The list price of an iPod speaker is $67.50. It is on sale for 20% off. and there is a 6% sale tax. The store calculates the sale tax on the $67.50. and then takes 20% off the total. Jonathan wants them to take the discount first and then calculate the tax. Reasoning that the tax will be calculate based on the lower price and he will pay less.Is Jonathan correct? How much will he pay each way?
Answer by stanbon(75887)

The list price of an iPod speaker is $67.50. It is on sale for 20% off. and there is a 6% sale tax. The store calculates the sale tax on the $67.50. and then takes 20% off the total. Jonathan wants them to take the discount first and then calculate the tax. Reasoning that the tax will be calculated based on the lower price and he will pay less. Is Jonathan correct? How much will he pay each way?
--------
Tax on 67.50 = 0.06*67.5 = $4.05 ; Total cost = 67.5+4.05 = $71.55
Tax if 20% discounted = 0.06*0.8(67.5) = $3.24; Total cost = $54+3.24 = $57.24
----
Jonathan is correct.
Cheers,
Stan H.
---------------